A manufacturer of shoes wants to entice you to a big contract. A pair of shoes that you like as a wholesaler is currently listed at $59. In order to make the trade the manufacturer prices the sneakers for you at $59 less 102 3 %, 5%, 1%, but only if you purchase 1000 or more. If you agree to this trade, then:

a) What is the net price factor? Remember: keep at least 6 decimal places if needed.

b) What does this NPF mean?

c) What is the net price per one pair?

d) What was the exact single rate of discount that was allowed? Provide answer as percent rounded to one decimal spot.
Answer #1

Given discount factors are 10%, 3%, 5%, 1%

a) Net price factor = (0.90)*(0.97)*(0.95)*(0.99) = 0.821056

b) The amount actually pays for the merchandise after the discount has been deducted.

c) Net price per one pair = $59 * 0.821056 = $48.4423

d) If he purchased 1000 or more the single rate of discount that was allowed was

1000*$59*(1-0.821056) = 10,557.696 = 10,557.7